If 6cm is added to each side of a square, its perimeter becomes 48cm. how long is each side of the square?
This is what I've done, but the answer seems wrong.
Let x be one of the sides of the square.
Then, 6 + x will be the rest of the sides of the square
4x + 6 + x = 48
(combine like terms)
5x + 6 = 48
(subtract 6 from both sides)
5x = 42
(divide 5 from both sides)
x = 8.4
I honestly don't know. Probably the "Let's" statement I've done wrong.

Answer by jim_thompson5910(35256)

Let x = original side of the square

"If 6cm is added to each side of a square", then the new side length of the square is x+6 cm

The perimeter is equal to 4 times the side length (since each side is equal and all sides are added up)

So P = 4s. But P = 48 and s = x+6, so

P = 4s

48 = 4(x+6)

From here, solve for x

48 = 4x+24

48-24 = 4x

24 = 4x

24/4 = x

6 = x

x = 6

So the side length of the original square is 6 cm.